A bungee cord stretches 2-4 times their unstretched length, and a jumper feels no more than 2.5-3.5g of acceleration. Assume the force in the cord has the math form kB=Jg3, where k and J are constants and B is the amount of stretch in the cord past the unstretched length, and g is gravitational acceleration 32.2. Design a cord (design the constants k and J) so the cord stretches 2.5 times the unstretched length, and the acceleration of the jumper doesn't exceed 3g, and the cord has zero stiffness at the bottom of a 400ft drop...

HELP :cc
Answer #1

let unstreched length be l and and spring constant be k hence 2.5l+l =400 hence l =114.29 ft and k*(400-114.29) =m(3g+g) g =32.2 ft/s^2 you have to provide jumper's mass to solve this question